A proton with a kinetic energy of 4.9×10−16 Jmoves perpendicular to a magnetic field of 0.26 T Part A What is the radius of its circular path Express your answer using two significant figures. 2 cIn Submit My Answers Give Up
Answer #1

mv2/r = qvB

m is the mass of the proton, v is the velocity, q is the charge and B is the applied magnetic field

r = mv/qB

0.5mv2 = 4.9 \times 10-16

v2 = 4.9 \times 10-16/ (0.5 \times 1.67 \times 10-27)

v = 0.766045 \times 106 m/s

So, r = mv/qB

r = 1.67 \times 10-27\times0.766045 \times 106 / ( 1.6 \times 10-19\times 0.26)

r = 0.030752 m